

********************************************************************
UARS WEEKLY STATUS REPORT - 02 JUL 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM was turned on as scheduled on June 30 after being off for 6 days.
Completion of the turn-on procedure was delayed by telemetry problems at
White Sands, resulting in the loss of solar data on June 30. Normal
operations, including the planned "monthly" calibration scans of
deuterium lamp spectra, were resumed on July 1.
We are evaluating the recently reprocessed V20 products before their
release to the DAAC.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM/AXIS activated on June 30. While AXIS is active, it produces
energy deposition from 16 pixels based on precipitating electrons derived
from measured X-ray fluxes.
The PEM spectrometers continue to be active under PMON control. While
under PMON control, the PEM instrument computed energy deposition from its
particle spectrometers.
The FOT and the PEM team are completing close-out documentation on
the HEPS Energy Processor trouble; however progress has been hampered
with the departure of Peter Thomas.
***********
HALOE
***********
HALOE continues successful operations with all engineering parameters
within nominal operating range. HALOE science data products continue to be
of excellent quality.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
10 days CLA (7.0) PROD (L1-L3) (475R-71C-8F=396L)
0 days GPI (4.05) PROD (MLS V5 3AL) (39R-0C-1F=39L)
50 days WIN (5.11) PROD (L1-L3) (1540R-837C-5F=698L)
***********
FOT
***********
REPORTING PERIOD
This report covers 19 June 1999 (Orbit 42459 , GMT Day 170) through 25 June
(GMT Day 176).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
ACRIM 173/2334 42518 Off
SUSIM 173/2334 42518 Off
MLS 174/0129 42519 On
The beta angle ranged between the angles of 22.7 degrees to 33.5 degrees
for this report period. The beta angle is now increasing toward a maximum
angle of 34 degrees on 26 June 1999 (DOY 177).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.1 degrees C. Additional battery performance data
for each battery on 18 June 1999 and 25 June 1999 are provided in Appendix
D.
The clock error ranged from 9.9 to -10.5 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 04
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% MLS 08
% FOT 02
% ACRIM 01
% SUSIM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 01
No Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/130 - 2287.496913
Transponder B: 99/151 - 2287.499087
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/119 show the observatory data loss to be
24 hours, 13 minutes, 08.340 seconds (an increase of 0.000 seconds since
the last report period). This is a 0.0356 percent data loss which equals a
99.9644 percent data capture for the mission.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT personnel are supporting the IMOC (SCS-21) development effort. The
current work is a continuation of the SCS-21 UARS implementation effort
done last year which culminated in a level 1 demonstration.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT
The end-to-end testing has been delayed until a new FORMATS software
release can be made to accommodate the special Y2K planning aids.
MEETINGS
None
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver (reverse-to-forward) is scheduled for 10
July 1999 (DOY 191).
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 25 June 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
42497 172/1355 99-207 OBC Clk Adj 0.00 to 20.6 msec/day
42517 173/2149 99-208 OBC Clk Adj 20.6 to 0.00 msec/day
42518 173/2334 99-209 ACRIM Power OFF
42518 173/2334 99-210 SUSIM Power OFF
42519 174/0129 99-211 MLS Power ON
42522 174/0721 99-212 MLS SCAN ACT OFF
42529 174/1700 99-213 MLS SCAN ACT ON
42534 175/0136 99-214 MLS MLS Stop/Run
42537 175/0710 99-215 MLS MLS Stop/Run
42543 175/1556 99-216 MLS MLS Stop/Run
42547 175/2231 99-217 MLS SCAN ACT OFF
42548 176/0014 99-218 MLS SCAN ACT ON/MPL Load
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 25 June 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSPBIMON Y-HI PTS P/B Cur Mon 42489-90,94
HRPSEQBIMON Y-HI HRDI Q/B Cur Mon 42521
MLSCANACTTMP Y-LO MLS Scan Act Temp 42522
MLSCANACTTMP Y-LO MLS Scan Act Temp 42541
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 25 June 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
42497 172/13:55 (21 June 99) from 00.0 to 20.6 msec/day
42517 173/21:56 (22 June 99) from 20.6 to 00.0 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 25 June 1999
18 June 1999 - GMT Day 169 Beta = 22.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -95.2/-123.2 -2.73/-3.99 +0.0/ +0.0 16.5 0.000 0.0
2 +56.0/-5.6 +7.09/+5.15 +30.4/-15.2 23.8 1.021 22.7
3 +22.4/-50.4 +5.79/+4.19 +29.6/-16.0 23.7 1.024 22.5
25 June 1999 - GMT Day 176 Beta = 33.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -95.2/-123.2 -2.42/-3.68 +0.0/ +0.0 16.5 0.000 0.0
2 +22.4/-39.2 +6.44/+5.47 +29.2/-15.2 24.2 1.023 22.5
3 +22.4/-67.2 +5.47/+4.51 +28.8/-15.6 24.2 1.025 22.2
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 25 June 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-028 FOT 42493 172/0853 14170 Generic Late Acq.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-028 FOT see list June 99 14170 Generic Late Acq.
SEVNTFW99/26